1)    You work for Bellevue Window Products. While

performing an analysis for a new window product,

you found a report from last year that provided

the following information regarding the

manufacture of a similar product: annual production

rate = 40,000 units; selling price = $70 per

unit; fixed production cost = $240,000 per year;

variable production cost = $1,700,000 per year;

variable selling expenses = $96,000 per year. As

a first-cut, you decide to use this information to

estimate (a) the breakeven production rate per

year, (b) the company’s profit last year, and (c)

the annual production rate that would generate a

profit of $1,000,000 per year. What are your

estimates?



1
Expert's answer
2021-09-19T18:35:12-0400

(a) the breakeven production rate pery ear is: BEP = 240,000/(70 - 1,796,000/40,000) = 9,561.75 units.

(b) The company’s profit last year is:

TP = 70×40,000 - (1,796,000 + 240,000) = 764,000.

(c) The annual production rate that would generate a profit of $1,000,000 per year is above 40,000 units.